TMS Meeting Minutes, January 13, 2022

Attendees: Andrew, Christian, Heejoo, John, Olga, Trenton, Yang (Zoom Meeting)

Update on TMS Shutter Testing:
  • Heejoo has been running the TMS shutter tests during the daytime for 2-3 hours at a time
  • Has confirmed that if shutter is left open, there are no issues, even when running for 5 hours
  • Mark W. from Etalon suggested the test run today, which confirmed that the error belongs only to shutter operation
  • We have one last USB cable left to test with the shutter control
  • LeRoy to swap the cable soon and Heejoo will complete this test soon
  • Heejoo updated team on the details of the cables used in the test
  • 3021001-03, U023-003, U021-003 cables from Digi-Key Electronics
  • Mark W. from Etalon suggested that there could be an issue with TCP/IP for shutter control, if it's not the USB cable
  • If it's not any of the things we have identified as possible sources of error thus far, then we may need to have an Etalon technician come out to solve the problem
  • Even if we don't solve this right now, we can still continue work on commissioning TMS, just without the use of the shutter
Commissioning:
  • Unfortunately, we don't have a time slot to test TMS tonight
  • The earliest time when LBC might be on is tomorrow night, but it seems this might be out of the question, because there are other engineering needs
  • Starting Sat. 1/22, OSURC will very likely be using LBC's to take data for monitoring of galaxies
  • There are engineering nights before the OSURC run, but those are likely to be over-subscribed, as the left front rotator should be repaired by then and LUCI AO needs to be checked out, as well as MODS
  • These are full engineering nights, though, so we may have a chance to use the LBC's, but the highest priority for the first engineering night is LUCI AO checkout (though this is best conducted with the highest seeing)
  • After Heejoo and Trenton stabilize the collimator mounts, Andrew is most interested in taking a reference baseline with TMS and observing the amount of correction required with FPIA relative to TMS
Mountain Trip:
  • Trenton (possibly Heejoo) planning to head up the mountain to stabilize mounts and install the MPO fiber casette
  • Heejoo to put together a tasklist for the trip
Thermal Analysis of Honeycomb Mirrors:
  • Trenton making progress on thermal modelling with Brandon Challifoux's help
  • John noted that the hardpoint stiffnesses are ~100 N/m
  • Heejoo tasked Trenton with preparing a set of time-oriented goals to help motivate a timeline for the project
TMS Logging Discussion:
  • Heejoo concerned about the growing size of the log files
  • Yang plans to revise the TMS logging algorithms to reduce the footprint on the drive
  • Planning to only copy the files explicitly listed on the Wiki page
  • Olga wondering if it might be better to place the log files in /lbt/logs
  • John suggests keeping the data on robs for now, but once we have the official TMS software up and running, the file logging will go to the database, not robs
ESO Code:
  • Ron at ESO has a code (Sensitizer) which uses 5th order aberration theory to produce PSF's and obtain FWHM etc. and has a source extractor
  • Might be possible for Ron to do some consulting and improve on active optics at LBT
  • Ron's code uses a least squares best fit to the optical model with perturbations applied to parameters of the model (mirror shape, position, detector shape, position, etc.)

-- TrentonBrendel - 13 Jan 2022
Topic revision: r1 - 13 Jan 2022, TrentonBrendel
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback